f(x)=-x^2-4x
intervals f is increasing and decreasing.
Give the equation of the line that represents the axis of the parabola.

Answer by nerdybill(7384)

f(x)=-x^2-4x
intervals f is increasing and decreasing.
Give the equation of the line that represents the axis of the parabola.
.
axis of parabola:
x = -b/(2a)
x = -(-4)/(2(-1))
x = 4/(-2)
x = -2 (axis of parabola)
.
Since the coefficient associated with the x^2 term is negative, it is a parabola that opens downwards.
Therefore:
f is increasing in the region:
(-oo, -2)
and
f is decreasing in the region:
(-2, +oo)
where
oo is for infinity